Assume 2 𝑙𝑏 of gas, for which 𝑅 = 26 𝑓𝑑.𝑙𝑏 𝑙𝑏.°𝑅 and π‘˜ = 1.10, undergoing a polytropic process from 𝑝1 = 15 π‘π‘ π‘–π‘Ž and 𝑑1 = 100℉ to 𝑝2 = 75 π‘π‘ π‘–π‘Ž and 𝑉2 = 3.72 𝑓𝑑3 . For both nonflow and steady-flow (βˆ†π‘ƒ = 0, βˆ†πΎ = 0) processes, determine (a) 𝑛 in 𝑝𝑉 𝑛 = 𝐢; (b) βˆ†π», βˆ†π‘ˆ, π‘Žπ‘›π‘‘ βˆ†π‘†; (c) ∫ 𝑝𝑑𝑉 π‘Žπ‘›π‘‘ βˆ’ ∫ 𝑉𝑑𝑝; (d) π‘Š π‘Žπ‘›π‘‘ 𝑄.


1
Expert's answer
2021-10-31T18:12:09-0400

"m= 2 \;lb \\

R = 26 \;ft \;lb /lb \;°R \\

k = 1.1 \\

p_1 = 15 \;psia = 15 \times 144 = 2160 \; lbf/ft^2\\

T_1 = 100 \;°F = 559.67 \;°R \\

p_2 = 75 \;psia = 75 \times 144 = 10800 \;lbf/ft^2\\

V_2 = 3.72 \;ft^3"

Using ideal gas equation:

"p_1V_1 = mRT_1 \\

V_1 = \frac{mRT_1}{p_1} \\

= \frac{2 \times 26 \times 559.67}{2160} \\

= 13.48 \;ft^3

p_2V_2 =mRT_2 \\

T_2 = \frac{p_2V_2}{mR} \\

= \frac{10800 \times 3.72}{2 \times 26} \\

= 772.62 \; °R"

Using relations

"\frac{c_p}{c_v} = k \\

c_p -c_v = R \\

c_p = \frac{kR}{k-1} \\

c_p = \frac{1.1 \times 26}{1.1-1} \\

c_p = 286 \; ft \; lbf/lb \;°R \\

c_v = \frac{R}{k-1} \\

c_v = \frac{28}{1.1-1} \\

c_v = 260 \; ft \;lbf/lb \; °R"

(a)

"p_1V_1^n = p_2V_2^n \\

\frac{p_1}{p_2} = (\frac{V_2}{V_1})^n \\

ln(\frac{p_1}{p_2}) = ln(\frac{V_2}{V_1})^n \\

n = \frac{ln(\frac{p_1}{p_2})}{ln(\frac{V_2}{V_1})} \\

n = \frac{ln(\frac{15}{75})}{ln(\frac{3.72}{13.48})} \\

n=1.25"

(b)

"ΔH = H_2 -H_1 =m(h_2-h_1) =m(u_2p_2V_2 -u_1 -p_1V_1) \\

ΔH = m(u_2-u_1 +p_2V_2 -p_1V_1) \\

=mc_v(T_2-T_1) + p_2V_2 -p_1V_1 \\

= 2 \times 260(772.62-559.67) + 10800 \times 3.72 – 2160 \times 13.48 \\

= 110734+40176 -29116.8 \\

= 121793.2 \;lbf \cdot ft \\

1 \;BTU = 778 \;lbf \cdot ft \\

ΔH = \frac{1211793.2}{778}=156.5 \;BTU \\

ΔU = mc_v(T_2-T_1) = 2 \times 260 (772.62 -559.67) \\

= 110734 \;lbf \cdot ft \\

ΔU = \frac{110734}{778} = 142.3 \;BTU \\

ΔS = mc_p ln(\frac{V_2}{V_1}) + mc_v ln(\frac{p_2}{p_1}) \\

= 2 \times 286 \times ln(\frac{3.72}{13.48}) + 2 \times 260 \times ln(\frac{75}{15}) \\

= 100.47 \;lbf \cdot ft/ °R \\

ΔS = \frac{100.47}{778} = 0.129 \;BTU/°R"

(c)

"pV^n = c \\

p = \frac{c}{v^n} \\

\int^2_1 \frac{c}{v^n} dV = c \int^2_1 V^{-n} dV = c[\frac{V^{1-n}}{1-n}]^2_1 \\

= c[\frac{V_2^{1-n} -V_1^{1-n}}{1-n}] \\

= \frac{cV_2^{1-n} -cv_1^{1-n}}{1-n} \\

= \frac{p_2v_2^nv_2^{1-n} -p_1v_1^nV_1^{1-n}}{1-n} \\

= \frac{p_1V_1 -p_2V_2}{n-1} \\

= \frac{2160 \times 13.48 -10800 \times 3.72}{1.25-1} \\

= -44236.8 \; lbf \cdot ft \\

\int pdv = \frac{-44236.8}{778} = -56.86 \;BTU = -57 \;BTU \\

-\int Vdp = -\int^2_1 (\frac{c}{p})^{1/n} dp \\

= -c^{1/n} \int^2_1 p^{-1/n} dp \\

= -c^{1/n} [\frac{p^{1- 1/n}}{1- 1/n}]^2_1 \\

= -c^{1/n} [\frac{p_2^{1 – 1/n} -p_1^{1 -1/n}}{1 -1/n}] \\

= \frac{-c^{1/n} p_2^{1-1/n} +c^{1/n}p_1^{1-1/n}}{1-1/n} \\

= \frac{-p_2^{1/n}v_2p_2^{1-1/n} +p_1^{1/n}v_1p_1^{1-1/n}}{1-1/n} \\

= \frac{n(p_1V_1 -p_2V_2)}{n-1} \\

= 1.25(\frac{p_1V_1 -p_2V_2}{n-1}) \\

= 1.25 \int^2_1 pdV \\

= 1.25 \times -57 \\

= -71.25 \;BTU \\

-\int Vdp = -71.25 \;BTU"

(d) For a polytropic process:

For non-flow process

"W_{1-2} = \int pdV = -57 \;BTU"

For steady flow

"W_{1-2} = -\int Vdp = -71.3 \;BTU"

From the first law of thermodynamics

"Q=W+ΔU \\

= -57 +142.3 \\

= 85.3 \;BTU"